# Entropy and type-token ratio in gigaword corpora

### v2

@@ I Introduction
Nonetheless, the production of natural language texts is constrained by statistical laws. These empirical rules highlight patterns that occur in word frequency and vocabulary growth across languages. On the one hand, Zipf law [23] states that the distribution of word frequencies in a text follows a power law. On the other hand, Heaps law [24, 25] dictates that the number of types in a text increases sublinearly with its total number of tokens. These laws, among others, are found in numerous linguistic corpora [26, 27, 28] and help linguists and computational scientists model language, offering insights into communication efficiency and the structure of human language systems.
@@ II Data
The preprocessing criteria of the texts forming the corpora follow those established by Gerlach et al. [37], including tokenization, removal of punctuation and numbers, conversion to lowercase, and preservation of accents. The corpora which are publicly available may be found in [38]. Below is a brief overview, as summarized in Table 2.
@@
The second corpus (SPA) is part of the Corpus del Español [45], constituing a large collection of texts in Spanish from web pages of 21 Spanish-speaking countries, gathered between 2013 and 2014. After preprocessing, this corpus comprises 1,887,720,265 tokens and 3,699,411 types, of which 1,900,485 (51%) are hapax legomena. The Corpus del Español is widely used in quantitative linguistics of Spanish, from multi-dimensional analysis of register variation [46] to comparisons between corpora and live expositions to certain language forms [47].
The third corpus, the Turkish subcorpus of CC100 (TRCC100), is a collection of web crawl data from Common Crawl [48] in Turkish compiled by Facebook AI team using CCNet [49, 50]. After preprocessing, this corpus counts with 2,640,929,269 tokens and 9,804,688 types, and 5,159,681 hapax legomena (53% of the types).
@@ III.2 Asymptotic expression
$H=\frac{a}{K_{V,a}}\sum\limits_{r=1}^{V}\frac{\log r}{r^{a}}+\log K_{V,a},$ (5)
@@
$\sum\limits_{r=1}^{V}\frac{\log r}{r}=\gamma_{1}-\gamma_{1}(V+1),$ (8)
@@
$\sum\limits_{r=1}^{V}\frac{\log r}{r}\sim\frac{\left(\log V\right)^{2}}{2}$ (9)
for large $V$ (see leading order of Eq. (1.12) in [78] or Eq. (3.1) of [77]). Compiling the above relations, one arrives at [79, 80, 81]
$H\sim\frac{1}{2}\log V+\log\log V.$ (10)
@@
$H_{p}(r)=-\sum\limits_{r^{\prime}=1}^{r}p_{r^{\prime}}\log p_{r^{\prime}},$ (11)
@@
$H\sim\frac{\beta}{2}\log L+\log\log L,$ (12)
where the dependence on $\alpha$ is negligible in the limit of large sample sizes. Even though the first term of Eq. (12) dominates over the second term for $L\gg 1$, it is relevant to note that our datasets involve $L\sim 10^{9}$ and therefore the contribution of the second term can not be neglected.
@@ III.3 Validation
$H_{\text{fit}}\left(L\right)=p_{1}\left(\frac{\beta}{2}\log L+\log\log L\right)+p_{2}$ (13)
where $p_{1}$ and $p_{2}$ are free fitting parameters, independent of $L$, for the six datasets, and the $\beta$ values are the ones obteined in Fig. 1.
The results are shown with dashed lines in Fig. 2 and numerically in Table 3, where we also present a few metrics to evaluate the goodness of the fit for Eq. (13) to the data, namely, the coefficient of determination, $\rho^{2}$, the Spearman correlation coefficient, $\rho_{\mathrm{s}}$, the distance correlation coefficient $\rho_{\mathrm{d}}$ (a measure suitable for nonlinear dependences [82, 83]), and the $p$-value. We see a clear agreement between data and the fit of Eq. (13), except for the TWTR corpus due to its poor statistics, as mentioned above. Specifically, the worst values of the metrics in Table 3 (without considering the TwTR corpus) are $\rho^{2}_{\mathrm{min}}=0.91$, $r_{\mathrm{s,min}}=0.93$, and $\rho_{\mathrm{d,min}}=0.90$.
@@ IV Type-token ratio
It is worth notint that there are multiple methods for calculating the type-token ratio. Throughout this work, we employ an over-all approach, which is explicitly influenced by text length. Alternative metrics, such as mean-segmental approaches, are designed to avoid scaling with sample size . It is important to note that if we had used mean-segmental methods, our results would have been significantly different. This point is discussed further in Section VI.
@@ V Entropy and type-token ratio
$H$ $\sim\frac{\beta}{2\left(1-\beta\right)}\log TTR^{-1}+\log\log TTR^{-1},$ (15)
@@
$H_{\text{fit}}\left(TTR\right)$ $=p_{3}\left[\frac{\beta}{2\left(\beta-1\right)}\log TTR^{-1}\right.$ (16)
$\left.+\log\log TTR^{-1}\right]+p_{4},$
to the data, $p_{3}$ and $p_{4}$ being fitting parameters. Using the values for $\beta$ extracted from the fits of Fig. 1, we obtain the fits shown in Fig. 6 (dashed lines) and in Table 5 using the plug-in estimator. These results show an overall good agreement between the asymptotic expression in Eq. (16) and the data. Equation (16) inherits the validity for large $L$ from Eq. (12), and hence we observe an asymptotic validity of the relation in Eq. (15). This fit clarifies the relationship between $H$ and $TTR$ for small values of $TTR$, which is based on logarithmic functions. This relationship is expressed by Eq. (15) and depends on the validity of Zipf and Heaps laws. Similar results using the NSB estimator are shown in Appendix D.
@@ VI Conclusions
A word of caution is here in order. Our results depend on the over-all approach used when computing the type-token ratio. Using instead a mean-segmental approach would require fixing a specific text length [1]. This adjustment would alter the results presented in Section V and invalidate the use of the variable $L$ for analytically linking entropy and the type-token ratio. Then, a positive correlation would be found between the mean-segmental (or standardized) type-token ratio and entropy, as both metrics reflect lexical richness. In this study, we do not obtain this positive correlation because we use the over-all approach to compute the type-token ratio.
@@
Finally, it would be interesting to investigate in future works if the functional relation found here between two diversity measures also holds in datasets from different fields such as ecology and genetics, where Zipf and Heaps law are observed [93, 94]. Ultimately, the connection between entropy and richness could be traced back to information efficiency in complex systems.
@@ Acknowledgements.
We thank M. A. Tugores for insightful discussions about the corpora analysis, and J. De Gregorio and E. G. Altmann for valuable comments and references. This work was partially supported by the Spanish State Research Agency (MICIU/AEI/10.13039/501100011033) and FEDER (UE) under project APASOS (PID2021-122256NB-C21) and the María de Maeztu project CEX2021-001164-M. The authors acknowledge the use of AI tools to assist with the writing of the manuscript.
@@ Appendix A Computation of the Entropy and TTR
Given a compiled corpus of length $L_{\text{T}}$, we consider fragments with lengths $l\in[2.5\times 10^{6},L_{\text{T}}]$ in steps of $\Delta l=2.5\times 10^{6}$ tokens. For each value of $l$, we take 25 random fragments of $l$ tokens from the total of $L_{\text{T}}$ tokens and compute the distributions of $H$, $V$, and $TTR$. The codes used for the analysis of the corpora are available in [38].
@@ Appendix B Contribution to the entropy of the second regime of the frequency distribution
One option is to use ratio of frequencies as a proxy to determine the contribution to the entropy of the words ranked beyond the critical value, i.e., the rank at which the Zipf exponent changes from around 1 to approximately 2. As shown in [95], the contribution of the ranks below the critical value ($b$ in [95], $r_{c}$ in Table 6) is around 80%.
@@
$R_{H}$ $=\frac{\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\log p_{r_{i}}}{\sum\limits_{r_{i}=1}^{V}p_{r_{i}}\log p_{r_{i}}}$
$=\frac{\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\log p_{r_{i}}}{\sum\limits_{r_{i}=1}^{r_{c}-1}p_{r_{i}}\log p_{r_{i}}+\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\log p_{r_{i}}},$ (17)
@@
$R^{\prime}_{H}=\frac{c_{2}\sum\limits_{r_{i}=r_{c}}^{V}r_{i}^{-a_{2}}\log\left[\frac{c_{2}}{L}r_{i}^{-a_{2}}\right]}{c_{1}\sum\limits_{r_{i}=1}^{r_{c}-1}r_{i}^{-a_{1}}\log\left[\frac{c_{1}}{L}r_{i}^{-a_{1}}\right]+c_{2}\sum\limits_{r_{i}=r_{c}}^{V}r_{i}^{-a_{2}}\log\left[\frac{c_{2}}{L}r_{i}^{-a_{2}}\right]}$ (21)
@@ Appendix C Range selection for the fits
As stated in Section III, to select the appropriate range in which to perform the fit of Eq. 13 we search the values in which the data that we are fitting is not affected either by the noisy small $V$ regime or by the artificially low $\sigma_{H}$ observed when the sample size approaches the corpus size. For example, the range for SPA corpus is $0.0250.0025<\sigma_{H}<$ 0.025. This heuristic $\sigma_{H}$ condition works well for all the corpora, as seen in Fig. 8, with the exception of the Turkish ones. In both Turkish corpora, $\sigma_{H}$ is systematically higher than in the Spanish and English corpora, and has a different shape. Hence, in the case of Turkish corpora we choose a range of high vocabularies far from the beginning and the end of the sampling range for the TRCC100, corresponding to $902249,370<V<5,500,$ 902, and all the available data for the TwTR.
@@ Appendix D Measurements using NSB estimator
Figure 10: $H$ in terms of $TTR$ for the SPGC (a), SPA (b), TRCC100 (c), TwEN (d), TwES (e), and TwTR (f) corpora for $L>2.5\cdot 10^{6}$ tokens using the NSB estimator. With a dashed line we plot the fit to Eq. (16), whose goodness is assessed in Table 8. Inside the insets we show the data and the fit in the confidence interval, as exemplified in Fig. 4.

### v3

@@ I Introduction
Nonetheless, the production of natural language texts is constrained by statistical laws. These empirical rules highlight patterns that occur in word frequency and vocabulary growth across languages. On the one hand, Zipf law [23] states that the distribution of word frequencies in a text follows a power law. On the other hand, Heaps law [24, 25] dictates that the number of types in a text increases sublinearly with its total number of tokens. These laws, among others, are found in numerous linguistic corpora [26, 27, 28] and help linguists and computational scientists model language, offering insights into communication efficiency and the structure of human language systems. Their persistence among different multilingual corpora is rooted in the balance between repetition and novelty in human communication: a few common words are used very frequently for efficiency, while rarer words carry specificity and keep appearing as texts grow, though at a decreasing rate. Several generative models have been proposed to explain Zipf and Heaps laws and other statistical features of natural language, often highlighting their links to cognitive processes [29, 30, 31].
@@ II Data
The preprocessing criteria of the texts forming the corpora follow those established by Gerlach et al. [40], including tokenization, removal of punctuation and numbers, conversion to lowercase, and preservation of accents. The corpora which are publicly available may be found in Ref. [41]. Below is a brief overview, as summarized in Table 2.
@@
The second corpus (SPA) is part of the Corpus del Español [48], constituting a large collection of texts in Spanish from web pages of 21 Spanish-speaking countries, gathered between 2013 and 2014. After preprocessing, this corpus comprises 1,887,720,265 tokens and 3,699,411 types, of which 1,900,485 (51%) are hapax legomena. The Corpus del Español is widely used in quantitative linguistics of Spanish, from multi-dimensional analysis of register variation [49] to comparisons between corpora and live expositions to certain language forms [50].
The third corpus, the Turkish subcorpus of CC100 (TRCC100), is a collection of web crawl data from Common Crawl [51] in Turkish compiled by Facebook AI team using CCNet [52, 53]. After preprocessing, this corpus contains 2,640,929,269 tokens and 9,804,688 types, and 5,159,681 hapax legomena (53% of the types).
@@ III.2 Asymptotic expression
$H=\frac{a}{K_{V,a}}\sum\limits_{r=1}^{V}\frac{\ln r}{r^{a}}+\ln K_{V,a},$ (5)
@@
$\sum\limits_{r=1}^{V}\frac{\ln r}{r}=\gamma_{1}-\gamma_{1}(V+1),$ (8)
@@
$\sum\limits_{r=1}^{V}\frac{\ln r}{r}\sim\frac{\left(\ln V\right)^{2}}{2}$ (9)
for large $V$ (see leading order of Eq. (1.12) in Ref. [81] or Eq. (3.1) of [80]). Compiling the above relations, one arrives at [82, 83, 84]
$H\sim\frac{1}{2}\ln V+\ln\ln V.$ (10)
Throughout the manuscript, we use the symbol $\sim$ to denote that two functions are asymptotically equivalent. Specifically, if
$\lim_{n\to\infty}\frac{f(n)}{g(n)}=1,$ (11)
we write $f\sim g$ as $n\to\infty$.
@@
$H_{p}(r)=-\sum\limits_{r^{\prime}=1}^{r}p_{r^{\prime}}\ln p_{r^{\prime}},$ (12)
@@
$H\sim\frac{\beta}{2}\ln L+\ln\ln L,$ (13)
where the dependence on $\alpha$ is negligible in the limit of large sample sizes. Even though the first term of Eq. (13) dominates over the second term for $L\gg 1$, it is relevant to note that our datasets involve text lengths on the order of $10^{9}$ tokens and therefore the contribution of the second term can not be neglected.
@@ III.3 Validation
$H_{\text{fit}}\left(L\right)=p_{1}\left(\frac{\beta}{2}\ln L+\ln\ln L\right)+p_{2}$ (14)
where $p_{1}$ and $p_{2}$ are fitting parameters, independent of $L$, for the six datasets, in order to align the asymptotic approximation to the empirical data. In other words, Eq. (14) is a semi-empirical approach that retains the same functional behaviour of $H(L)$ for all corpora with $p_{1}$ and $p_{2}$ being specific for each dataset. This precisely follows the spirit of other statistical laws [26]. Also, the $\beta$ values are the ones obtained in Fig. 1.
The results of a nonlinear least squares (NLS) fitting [85] are shown with dashed lines in Fig. 2 and numerically in Table 3, where we also present a few metrics to evaluate the goodness of the fit for Eq. (14) to the data, namely, the coefficient of determination, $\rho^{2}$, the Spearman correlation coefficient, $\rho_{\mathrm{s}}$, the distance correlation coefficient $\rho_{\mathrm{d}}$ (a measure suitable for nonlinear dependences [86, 87]), and the $p$-value. We see a clear agreement between data and the fit of Eq. (14), except for the TWTR corpus due to its poor statistics, as mentioned above. Specifically, the worst values of the metrics in Table 3 (without considering the TwTR corpus) are $\rho^{2}_{\mathrm{min}}=0.91$, $r_{\mathrm{s,min}}=0.93$, and $\rho_{\mathrm{d,min}}=0.90$.
@@ IV Type-token ratio
It is worth noting that there are multiple methods for calculating the type-token ratio. Throughout this work, we employ an over-all approach, which is explicitly influenced by text length. Alternative metrics, such as mean-segmental approaches, are designed to avoid scaling with sample size . It is important to note that if we had used mean-segmental methods, our results would have been significantly different. This point is discussed further in Section VI.
@@ V Entropy and type-token ratio
$H\sim\frac{\beta}{2\left(1-\beta\right)}\ln TTR^{-1}+\ln\ln TTR^{-1},$ (16)
@@
$H_{\text{fit}}\left(TTR\right)$ $=p_{3}\left[\frac{\beta}{2\left(\beta-1\right)}\ln TTR^{-1}\right.$ (17)
$\left.+\ln\ln TTR^{-1}\right]+p_{4},$
to the data, $p_{3}$ and $p_{4}$ being fitting parameters. Using the values for $\beta$ extracted from the fits of Fig. 1, we obtain the NLS fits shown in Fig. 6 (dashed lines) and in Table 5 using the plug-in estimator. These results show an overall good agreement between the asymptotic expression in Eq. (17) and the data. Equation (17) inherits the validity for large $L$ from Eq. (13), and hence we observe an asymptotic validity of the relation in Eq. (16). This fit clarifies the relationship between $H$ and $TTR$ for small values of $TTR$, which is based on logarithmic functions. This relationship is expressed by Eq. (16) and depends on the validity of Zipf and Heaps laws. Similar results using the NSB estimator are shown in Appendix D.
@@ VI Conclusions
A word of caution is here in order. Our results depend on the over-all approach used when computing the type-token ratio. Using instead a mean-segmental approach would require fixing a specific text length [1]. This adjustment would alter the results presented in Section V and invalidate the use of the variable $L$ for analytically linking entropy and the type-token ratio. Then, since both metrics reflect lexical richness, a positive correlation would be found between the mean-segmental (or standardized) type-token ratio and entropy, as shown in the literature [91]. In this study, we do not obtain this positive correlation because we use the over-all approach to compute the type-token ratio.
@@
Importantly, we did not include punctuation marks in our analysis to stay consistent with the text-filtering approach used in [40]. However, this is an interesting direction for future work. Notably, Ref. [98] shows that punctuation marks behave statistically like common words and can even improve the fit to Zipfian distributions.
@@
Finally, it would be interesting to investigate in future work whether the functional relationship found here between two diversity measures also holds in datasets from other fields, such as ecology and genetics, where Zipfs and Heaps laws have been observed [99, 100]. Since our derivation shows that the entropy–TTR connection follows from these two laws, one would expect the same relationship to hold in any domain where they apply. However, our analysis has focused exclusively on language data, and testing this in other contexts could help assess the robustness and limitations of the connection.
@@ Acknowledgements.
We thank M. A. Tugores for insightful discussions about the corpora analysis, and J. De Gregorio and E. G. Altmann for valuable comments and references. This work was partially supported by the Spanish State Research Agency (MICIU/AEI/10.13039/501100011033) and FEDER (UE) under project APASOS (PID2021-122256NB-C21) and the María de Maeztu project CEX2021-001164-M.
@@ Appendix A Computation of the entropy and TTR
Given a compiled corpus of length $L_{\text{T}}$, we consider fragments with lengths $l\in[2.5\times 10^{6},L_{\text{T}}]$ in steps of $\Delta l=2.5\times 10^{6}$ tokens. For each value of $l$, we take 25 random fragments of $l$ tokens from the total of $L_{\text{T}}$ tokens and compute the distributions of $H$, $V$, and $TTR$. The codes used for the analysis of the corpora are available in Ref. [41].
@@ Appendix B Contribution to the entropy of the second regime of the frequency distribution
One option is to use ratio of frequencies as a proxy to determine the contribution to the entropy of the words ranked beyond the critical value, i.e., the rank at which the Zipf exponent changes from around 1 to approximately 2. As shown in Ref. [101], the contribution of the ranks below the critical value ($b$ in Ref. [101], $r_{c}$ in Table 6) is around 80%.
@@
$R_{H}$ $=\frac{\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\ln p_{r_{i}}}{\sum\limits_{r_{i}=1}^{V}p_{r_{i}}\ln p_{r_{i}}}$
$=\frac{\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\ln p_{r_{i}}}{\sum\limits_{r_{i}=1}^{r_{c}-1}p_{r_{i}}\ln p_{r_{i}}+\sum\limits_{r_{i}=r_{c}}^{V}p_{r_{i}}\ln p_{r_{i}}},$ (18)
@@
$R^{\prime}_{H}=\frac{c_{2}\sum\limits_{r_{i}=r_{c}}^{V}r_{i}^{-a_{2}}\ln\left[\frac{c_{2}}{L}r_{i}^{-a_{2}}\right]}{c_{1}\sum\limits_{r_{i}=1}^{r_{c}-1}r_{i}^{-a_{1}}\ln\left[\frac{c_{1}}{L}r_{i}^{-a_{1}}\right]+c_{2}\sum\limits_{r_{i}=r_{c}}^{V}r_{i}^{-a_{2}}\ln\left[\frac{c_{2}}{L}r_{i}^{-a_{2}}\right]}$ (22)
@@ Appendix C Range selection for the fits
As stated in Section III, to select the appropriate range in which to perform the fit of Eq. 14 we search the values in which the data that we are fitting is not affected either by the noisy small $V$ regime or by the artificially low $\sigma_{H}$ observed when the sample size approaches the corpus size. For example, the range for SPA corpus is $0.0250.0025<\sigma_{H}<$ 0.025. This heuristic $\sigma_{H}$ condition works well for all the corpora, as seen in Fig. 8, with the exception of the Turkish ones. In both Turkish corpora, $\sigma_{H}$ is systematically higher than in the Spanish and English corpora, and has a different shape. Hence, in the case of Turkish corpora we choose a range of high vocabularies far from the beginning and the end of the sampling range for the TRCC100, corresponding to 249,370$<V<$5,500,902, and all the available data for the TwTR.
@@ Appendix D Measurements using NSB estimator
Figure 10: $H$ in terms of $TTR$ for the SPGC (a), SPA (b), TRCC100 (c), TwEN (d), TwES (e), and TwTR (f) corpora for $L>2.5\times 10^{6}$ tokens using the NSB estimator. With a dashed line we plot the fit to Eq. (17), whose goodness is assessed in Table 8. Inside the insets we show the data and the fit in the confidence interval, as exemplified in Fig. 4.